# Model Inversion Attacks on Vision-Language Models: Do They Leak What They Learn?

### v1

@@ Abstract
Model inversion (MI) attacks pose significant privacy risks by reconstructing private training data from trained neural networks. While prior works have focused on conventional unimodal DNNs, the vulnerability of vision-language models (VLMs) remains underexplored. In this paper, we conduct the first study to understand VLMs’ vulnerability in leaking private visual training data. To tailored for VLMs’ token-based generative nature, we propose a suite of novel token-based and sequence-based model inversion strategies. Particularly, we propose Token-based Model Inversion (TMI), Convergent Token-based Model Inversion (TMI-C), Sequence-based Model Inversion (SMI), and Sequence-based Model Inversion with Adaptive Token Weighting (SMI-AW). Through extensive experiments and user study on three state-of-the-art VLMs and multiple datasets, we demonstrate, for the first time, that VLMs are susceptible to training data leakage. The experiments show that our proposed sequence-based methods, particularly SMI-AW combined with a logit-maximization loss based on vocabulary representation, can achieve competitive reconstruction and outperform token-based methods in attack accuracy and visual similarity. Importantly, human evaluation of the reconstructed images yields an attack accuracy of 75.31%, underscoring the severity of model inversion threats in VLMs. Notably we also demonstrate inversion attacks on the publicly released VLMs. Our study reveals the privacy vulnerability of VLMs as they become increasingly popular across many applications such as healthcare and finance.
@@ 1 Introduction
Figure 1: Overview of our proposed Model Inversion strategies for Vision-Language Models. Conventional MI typically targets unimodal DNNs, where the adversary seeks to reconstruct a training image $x=G(w)$ that maximizes the likelihood of a target class label $y$ under the target model $M_{DNN}$. The maximization is accomplished by repeating NN inversion steps to recover a high-fidelity reconstruction. In contrast, VLMs $M_{VLM}$ generate a sequence of tokens, and the target output $\mathbf{y}=(y_{1},y_{2},\dots,y_{m})$ is also a sequence of $m$ tokens. To address the unique nature of VLMs, we propose four MI strategies. (a) Token-based Model Inversion (TMI): We perform one gradient update to the latent variable $w$ after each generated token. This process continues for all $m$ tokens in the sequence, and the entire sequence-level inversion is repeated $K=N/m$ times. (b) Convergent Token-based Model Inversion (TMI-C): To ensure correctness of earlier tokens before generating subsequent ones, we propose updating $w$ for $K$ steps per token $y_{i}$, conditioning on the previous tokens $y_{<i}$. (c) Sequence-based Model Inversion (SMI): We compute a single gradient update to $w$ based on the average loss over all $m$ tokens, providing a global view of the sequence-level gradients. (d) Sequence-based Model Inversion with Adaptive Token Weighting (SMI-AW): We introduce adaptive token weights $\alpha_{i}$ for each token $y_{i}$ to dynamically emphasize tokens that could provide more essential feedback signals to guide the search toward a correct reconstructed image matching the target description.
Model Inversion (MI) attacks aim to reconstruct training data by exploiting information encoded within a trained model. These attacks pose significant privacy risks to unimodal deep neural networks (DNNs) (Fredrikson, Jha, and Ristenpart 2015; Zhang et al. 2020; Chen et al. 2021; An et al. 2022; Struppek et al. 2022; Kahla et al. 2022; Han et al. 2023; Nguyen et al. 2023b; Yuan et al. 2023; Nguyen et al. 2023a; Qiu et al. 2024; Peng et al. 2022; Ho et al. 2024; Koh et al. 2024; Struppek, Hintersdorf, and Kersting 2024). With the rapid advancement and widespread deployment of Vision-Language Models (VLMs) across various applications, an important and timely question arises: Are VLMs similarly vulnerable to Model Inversion attacks as unimodal DNNs? In this context, we define an MI attack as the task of reconstructing VLM’s training images by leveraging its textual input and output. Addressing this question is crucial for understanding and mitigating potential privacy threats in multimodal learning systems.
Research Gap. Existing MI attacks (Zhang et al. 2020; Chen et al. 2021; An et al. 2022; Struppek et al. 2022; Nguyen et al. 2023b; Qiu et al. 2024) primarily focus on unimodal DNNs, where the goal is to reconstruct private training images $x$ associated with a target label $y$. These methods typically pose inversion as an optimization problem that maximizes the likelihood of $y$ under the target model:
@@
Here, $M_{DNN}$ is a unimodal DNN trained on private data $\mathcal{D}_{priv}$, and $G$ represents a generative model, such as a Generative Adversarial Network (Goodfellow et al. 2014; Karras, Laine, and Aila 2019). The optimization is usually accomplished by performing $N$ inversion update steps to generate a reconstruction $x^{*}=G(w^{*})$ that approximates the original training sample in $\mathcal{D}_{priv}$ for a given label $y$.
@@
In this work, we introduce a suite of novel token-based and sequence-based model inversion strategies tailored for VLMs. Specifically, we propose four MI techniques (see Figure 1). Our token-based attacks leverage token-level gradients to optimize the reconstructed images. In contrast, our sequence-based attacks utilize gradients aggregated over the entire sequence, offering a global perspective for image optimization. Experimental results on three VLMs across three datasets demonstrate the effectiveness of our inversion attacks. Notably, human evaluation of the reconstructed images achieves an attack accuracy of 75.31%, highlighting the severity of model inversion threats in VLMs. Furthermore, we validate the generalizability of our approach on publicly available VLMs, reinforcing its practical applicability and security implications. Our key contributions are as follows:
  • We present a pioneering study of model inversion attacks on vision-language models, uncovering a novel security risk in the multimodal models.
  • We introduce a suite of novel inversion strategies tailored for VLMs, including two Token-based MI (TMI and TMI-C) and two Sequence-based MI attacks (SMI and SMI-AW).
  • The extensive experimental validation shows our proposed attacks, especially SMI-AW, achieve both high attack accuracy and good visual fidelity. Crucially, we showcase successful and high-fidelity inversion attacks against publicly available VLMs, underscoring the immediate and practical privacy risks posed by these models and the urgent need for robust defense mechanisms.
@@ 2 Problem Formulation
In this work, we present a pioneering study on model inversion attack in VLMs, which are increasingly adopted in real-world applications.
Threat Model. We consider a threat model where a VLM $M$ is pre-trained on broad data and fine-tuned on a private VQA dataset $\mathcal{D}_{priv}=\{(t,\mathbf{x},y)\}$, where $\mathbf{x}$ is the image, $t$ and $y$ are the textual input and correct textual answer.
For clarity, hereafter we use $M$ to denote a VLM and $M_{DNN}$ to refer to a unimodal DNNs. Using the tokenizer of $M$, the textual input $t$ and the textual answer $y$ are tokenized into sequences $\mathbf{t}=(t_{1},t_{2},\dots,t_{n})$ and $\mathbf{y}=(y_{1},y_{2},\dots,y_{m})$, respectively. We denote the full output sequence of $M$ given input $(\mathbf{t},\mathbf{x})$ as $M(\mathbf{t},\mathbf{x})$. The model’s prediction of the $i$-th token $y_{i}$, conditioned on the previous tokens $y_{<i}$, is denoted by $M(\mathbf{t},\mathbf{x},y_{<i})$.
Attacker’s Goal. Given a trained VLM $M$, the goal of a model inversion attack is to reconstruct a representative image $\mathbf{x}^{*}$ that reveals sensitive or private visual information from the private training image $\mathbf{x}$ in a data sample $(t,\mathbf{x},y)\in\mathcal{D}_{priv}$. Specifically, the adversary is given access to the trained model $M$, a textual input prompt $t$, and the corresponding target output $y$. The target is to synthesize an image $\mathbf{x}^{*}$ such that $M(\mathbf{t},\mathbf{x}^{*})=\mathbf{y}$ where $\mathbf{t}=(t_{1},t_{2},\dots,t_{n})$ and $\mathbf{y}=(y_{1},y_{2},\dots,y_{m})$ are the output token sequence associated with the input textual $t$ and textual output $y$. In other words, the model inversion attack seeks to infer a plausible visual input $\mathbf{x}^{*}$ that, when paired with the given input tokens $\mathbf{t}$, produces the high likelihood output sequence $\mathbf{y}$. This reconstructed image $\mathbf{x}^{*}$ is intended to approximate or reveal private features of the true image $\mathbf{x}$, thereby compromising the visual confidentiality of the training data.
Attacker’s Capabilities. We consider a white-box setting (Zhang et al. 2020; Chen et al. 2021; An et al. 2022; Struppek et al. 2022; Nguyen et al. 2023b; Qiu et al. 2024), where the attacker has full access to the VLM’s architecture, parameters, output responses (e.g., generated text or logits), input prompts $t$, and their corresponding ground-truth answers $y$. The attacker also has access to an auxiliary public dataset $\mathcal{D}_{pub}$ from the same domain as $\mathcal{D}_{priv}$.
@@ 3 Model Inversion Strategies for VLMs
Given a VLM $M$ trained on broad data and fine-tuned with a private VQA dataset $\mathcal{D}_{priv}=\{(t,\mathbf{x},y)\}$. Performing MI attacks directly in the image space is computationally expensive and often ineffective (Zhang et al. 2020). To reduce the search space of $x^{*}$, we follow conventional MI approaches for DNNs by leveraging a generative model $G$ trained on an auxiliary public dataset $\mathcal{D}_{pub}$ (Zhang et al. 2020; Chen et al. 2021; Nguyen et al. 2023b; Yuan et al. 2023; Struppek et al. 2022; Qiu et al. 2024). This allows us to shift the optimization from the high-dimensional image space to the lower-dimensional latent space of $G$, i.e., $x=G(w)$, where $w$ is the intermediate latent vector.
In contrast to conventional MI attacks targeting classification models-where the objective is to reconstruct an input image $x$ that yields a specific class label -VLMs generate token sequences, and the target output also represented as a sequence of tokens. This requires a reformulation of the MI objective to account for token generation. Our goal is to reconstruct a representative image $\mathbf{x}^{*}=G(w^{*})$ by optimizing the latent vector $w$ such that the generated image captures the semantic content of the private training image $\mathbf{x}$ that associates with description $y$.
In this section, we introduce four inversion strategies tailored for VLMs. The first two (TMI and TMI-C) are token-based approaches that leverage token-level gradients to optimize the reconstructed images. In contrast, the remaining two are sequence-based methods (SMI and SMI-AW) that aggregate gradients over the entire output sequence, providing a global perspective for inversion.
@@ 3.1 Token-based Model Inversion (TMI)
A natural approach is to treat the inversion process as a sequential update over individual token predictions. Given a target token sequence $\mathbf{y}$, we iteratively update the latent code $w$ after each generated token (see Figure 1 (a)). The details are in Algorithm 1.
@@
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},y_{2},\dots,y_{m}),N,\beta$
@@
4:Randomly initialize $w$.
5:for $k=1$ to $K$ do
6:  for $i=1$ to $m$ do
7:    $\hskip-48.36958pt\mathcal{L}=\mathcal{L}_{inv}(M(\mathbf{t},G(w),y_{<i}),y_{i})$ (2)
8:   $w=w-\beta\frac{\partial\mathcal{L}}{\partial w}$
@@
10:end for
Here, $N$ is the number of inversion steps, $\beta$ is the update rate of MI, $y_{<i}$ denotes the previous tokens. $\mathcal{L}_{inv}$ presents the inversion loss, guiding the generative model $G$ to produce images that induce the token $y_{i}$. We will discuss the design of $\mathcal{L}_{inv}$ in the next section.
The optimization is performed over multiple iterations, typically up to a update limit of $N$ inversion steps. At each iteration, each token contributes independently to the optimization process.
@@ 3.2 Convergent Token-based Model Inversion (TMI-C)
TMI performs a single update per token per iteration. However, VLMs generate each token $y_{i}$ based on the preceding tokens $y_{<i}$. To better align with this generative dependency, we propose Convergent Token-based Model Inversion (TMI-C), which updates the latent vector $w$ multiple times for each target token before proceeding to the next. Specifically, for each token $y_{i}$, we perform $K$ updates to $w$, thereby encouraging convergence of the token-level inversion subproblem before advancing to $y_{i+1}$ (see Figure 1 (b)). The details are presented in Algorithm 2.
@@
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},y_{2},\dots,y_{m}),N,\beta$
@@
4:Randomly initialize $w$.
5:for $i=1$ to $m$ do
6:  for $k=1$ to $K$ do
7:   Compute $\mathcal{L}$ using Eqn. (2).
8:   $w=w-\beta\frac{\partial\mathcal{L}}{\partial w}$
@@ 3.3 Sequence-based Model Inversion (SMI)
Token-based model inversion methods treat each token independently, optimizing the latent vector $w$ based on individual token-level losses. As the output of VLMs is a sequence of tokens, we propose Sequence-based Model Inversion (SMI), which performs a single gradient update to $w$ by averaging the loss across all $m$ tokens in the sequence (see Figure 1 (c)). By aggregating token-level losses into a unified objective, SMI leverages the interdependencies among tokens and provides more coherent gradients that reflects the structure of the full sequence. This global view encourages the model to recover a latent representation that is consistent across the entire sequence, rather than optimizing for each token in isolation. The details are presented in Algorithm 3.
@@
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},y_{2},\dots,y_{m}),N,\beta$
@@
3:Randomly initialize $w$.
4:for $k=1$ to NN do
5:   $\mathcal{L}=\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}_{inv}(M(\mathbf{t},G(w),y_{<i}),y_{i})$ (3)
6:  $w=w-\beta\frac{\partial\mathcal{L}}{\partial w}$

### v2

# Do Vision-Language Models Leak What They Learn? Adaptive Token-Weighted Model Inversion Attacks

@@ Abstract
Model inversion (MI) attacks pose significant privacy risks by reconstructing private training data from trained neural networks. While prior studies have primarily examined unimodal deep networks, the vulnerability of vision-language models (VLMs) remains largely unexplored. In this work, we present the first systematic study of MI attacks on VLMs to understand their susceptibility to leaking private visual training data. Our work makes two main contributions. First, tailored to the token-generative nature of VLMs, we introduce a suite of token-based and sequence-based model inversion strategies, providing a comprehensive analysis of VLMs’ vulnerability under different attack formulations. Second, based on the observation that tokens vary in their visual grounding, and hence their gradients differ in informativeness for image reconstruction, we propose Sequence-based Model Inversion with Adaptive Token Weighting (SMI-AW) as a novel MI for VLMs. SMI-AW dynamically reweights each token’s loss gradient according to its visual grounding, enabling the optimization to focus on visually informative tokens and more effectively guide the reconstruction of private images. Through extensive experiments and human evaluations on a range of state-of-the-art VLMs across multiple datasets, we show that VLMs are susceptible to training data leakage. Human evaluation of the reconstructed images yields an attack accuracy of 61.21%, underscoring the severity of these privacy risks. Notably, we demonstrate that publicly released VLMs are vulnerable to such attacks. Our study highlights the urgent need for privacy safeguards as VLMs become increasingly deployed in sensitive domains such as healthcare and finance. Code and additional experiments are provided in Supp.
@@ 1 Introduction
Figure 1: We conduct the first systematic study of MI attacks on VLMs. (A) Designed for the token-generative characteristics of VLMs, we introduce a set of token-level and sequence-level MI strategies to investigate VLMs’ privacy vulnerability (Sec. 3). Particularly, conventional MI typically targets unimodal DNNs, where the adversary seeks to reconstruct a training image $x=G(w)$ that maximizes the likelihood of a target class label $y$ under the target model $M_{DNN}$ by repeating NN inversion steps. In contrast, VLMs $M_{VLM}$ generate a sequence of tokens, and the target output $\mathbf{y}=(y_{1},\dots,y_{m})$ is also a sequence of $m$ tokens. To address the unique nature of VLMs, we introduce several MI strategies: Token-based Model Inversion (TMI), Convergent Token-based Model Inversion (TMI-C), and Sequence-based Model Inversion (SMI). (B) Building on the insight that output tokens differ in their degree of visual grounding, and hence their gradients vary in informativeness for reconstructing images during inversion, we propose Sequence-based Model Inversion with Adaptive Token Weighting (SMI-AW), a novel MI for VLMs (Sec. 4). SMI-AW adaptively adjusts each token’s gradient contribution according to its visual grounding, allowing the optimization to concentrate on visually grounded tokens and more effectively recover private training images. See Figure 2 for discussion of attention map analysis.
Model Inversion (MI) attacks aim to reconstruct training data by exploiting information encoded within a trained model. These attacks pose significant privacy risks to unimodal DNNs [fredrikson2015model, zhang2020secret, chen2021knowledge, an2022mirror, struppek2022plug, kahla2022label, han2023reinforcement, nguyen2023re, yuan2023pseudo, nguyen2023label, ho2024model, peng2024pseudo, qiu2024closer, tran2025random], The goal of MI attack is to reconstruct private training images $x$ associated with a target label $y$. These methods typically pose inversion as an optimization problem that maximizes the likelihood of $y$ under the target model:
@@
Here, $M_{DNN}$ is a unimodal DNN trained on private data $\mathcal{D}_{priv}$, and $G$ represents a generative model [goodfellow2014generative, adler2018banach, karras2019style]. The optimization is usually accomplished by performing $N$ inversion update steps to generate a reconstruction $x^{*}=G(w^{*})$ that approximates the training sample in $\mathcal{D}_{priv}$ for a given label $y$ (See Related Work in Supp).
Research Gap. With the rapid advancement and widespread deployment of Vision-Language Models (VLMs) across various applications [zhu2023minigpt, liu2024llavanext, team2024gemma, yao2024minicpm, wang2025internvl3, hong2025glm], an important and timely question arises: Are VLMs similarly vulnerable to Model Inversion attacks as unimodal DNNs? In this context, we define an MI attack as the task of reconstructing VLM’s training images by leveraging its textual input and output. Addressing this question is crucial for understanding potential privacy threats in multimodal systems.
@@
In this work, we conduct the first systematic investigation of MI attacks on modern VLMs (Figure 1). The token-generative nature of VLMs necessitates new MI attack designs beyond conventional unimodal approaches. To this end, we introduce a suite of token-based and sequence-based inversion strategies. Our token-based methods leverage token-level gradients to guide reconstruction, while our sequence-based methods aggregate gradients across the entire output sequence to provide a globally coherent optimization signal. Crucially, this framework reveals a key insight: output tokens differ substantially in their degree of visual grounding, and thus in how informative their gradients are for reconstructing images. Building on this observation, we propose Sequence-based Model Inversion with Adaptive Token Weighting (SMI-AW), which dynamically reweights token contributions using their visual attention strength, producing visually relevant gradients and enabling more accurate reconstruction of private training images.
We conduct experiments on a range of VLMs across multiple datasets to demonstrate the effectiveness of our inversion attacks. Notably, human evaluation of the reconstructed images achieves an attack accuracy of 61.21%, highlighting the severity of model inversion threats in VLMs. Furthermore, we validate the generalizability of our approach on publicly available VLMs, reinforcing its practical applicability and security implications. Our key contributions are as follows:
  • We present a pioneering study of MI attacks on VLMs, uncovering a security risk in the multimodal models.
  • We introduce a suite of inversion strategies tailored for token-generative nature of VLMs (Sec. 3).
  • Based on our observation that output tokens’ gradients differ in their informativeness for MI, we propose SMI-AW, which dynamically reweights token contributions in different inversion steps (Sec. 4).
  • The extensive experimental validation shows our proposed attacks, especially SMI-AW, achieve both good attack accuracy and good visual fidelity. Crucially, we showcase successful inversion attacks against publicly available VLMs, underscoring the immediate and practical privacy risks posed by these models (Sec. 5).
@@ 2 Problem Formulation
Threat Model. We consider a threat model where a VLM $M$ is trained on a private VQA dataset $\mathcal{D}_{priv}=\{(t,\mathbf{x},y)\}$, where $\mathbf{x}$ is the image, $t$ and $y$ are the textual input and correct textual answer. For clarity, hereafter we use $M$ to denote a VLM and $M_{DNN}$ to refer to a unimodal DNNs. Using the tokenizer of $M$, the textual input $t$ and the textual answer $y$ are tokenized into sequences $\mathbf{t}=(t_{1},t_{2},\dots,t_{n})$ and $\mathbf{y}=(y_{1},y_{2},\dots,y_{m})$, respectively. We denote the full output sequence of $M$ given input $(\mathbf{t},\mathbf{x})$ as $M(\mathbf{t},\mathbf{x})$. The model’s prediction of the $i$-th token $y_{i}$, conditioned on the previous tokens $y_{<i}$, is denoted by $M(\mathbf{t},\mathbf{x},y_{<i})$.
Attacker’s Goal. Given a trained VLM $M$, the goal of a model inversion attack is to reconstruct a representative image $\mathbf{x}^{*}$ that reveals sensitive or private visual information from the private training image $\mathbf{x}$ in a data sample $(t,\mathbf{x},y)\in\mathcal{D}_{priv}$. Specifically, the adversary is given access to the trained model $M$, a textual input prompt $t$, and the corresponding target output $y$. The attacker’s goal is to infer a plausible visual input $\mathbf{x}^{*}$ that produces the high likelihood output sequence $\mathbf{y}$ under the input tokens $\mathbf{t}$. This reconstructed image $\mathbf{x}^{*}$ is intended to approximate or reveal private features of the true image $\mathbf{x}$, thereby compromising the visual confidentiality of the training data.
Attacker’s Capabilities. We consider a white-box setting [zhang2020secret, chen2021knowledge, an2022mirror, struppek2022plug, nguyen2023re, qiu2024closer], where the attacker has full access to the VLM’s architecture, parameters, attention maps, output responses (e.g., generated text or logits).
@@ 3 Model Inversion Strategies for VLMs
We consider a VLM $M$ trained on a private VQA dataset $\mathcal{D}_{priv}=\{(t,\mathbf{x},y)\}$. Performing MI attacks directly in the image space is computationally expensive and often ineffective [zhang2020secret]. To reduce the search space of $x^{*}$, we follow conventional MI approaches for DNNs by leveraging a generative model $G$ trained on an auxiliary public dataset $\mathcal{D}_{pub}$ [zhang2020secret, chen2021knowledge, struppek2022plug, nguyen2023re, qiu2024closer]. This allows us to shift the optimization from the high-dimensional image space to the lower-dimensional latent space of $G$, i.e., $x=G(w)$, where $w$ is the intermediate latent vector.
In contrast to conventional MI attacks targeting classification models, where the objective is to reconstruct an input image $x$ that yields a specific class label, VLMs generate token sequences, and the target output also represented as a sequence of tokens. This requires a reformulation of the MI objective to account for token generation. In this section, we introduce new token-based and sequence-based MI methods. In Sec. 4, we further propose a novel MI with dynamic weighting to account for varying informativeness of different tokens’ gradients during inversion.
@@ 3.1 Token-based Model Inversion (TMI)
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},\dots,y_{m}),N,\lambda$
@@
4:for $k=1$ to $K$ do
5:  for $i=1$ to $m$ do
6:    $\hskip-45.52458pt\mathcal{L}=\mathcal{L}_{inv}(M(\mathbf{t},G(w),y_{<i}),y_{i})$ (2)
7:   $w=w-\lambda\frac{\partial\mathcal{L}}{\partial w}$
8:  end for
@@
A natural approach is to treat the inversion process as a sequential update over individual token predictions. Given a target token sequence $\mathbf{y}$, we iteratively update the latent code $w$ after each generated token (see Figure 1 (A) TMI). The details are in Algorithm 1. $N$ is the number of inversion steps, $\lambda$ is the update rate of MI, $y_{<i}$ denotes the previous tokens. $\mathcal{L}_{inv}$ presents the inversion loss, guiding the generative model $G$ to produce images that induce the token $y_{i}$. We discuss the design of $\mathcal{L}_{inv}$ in the Supp. The optimization is performed over multiple iterations, typically up to a update limit of $N$ inversion steps. At each iteration, each token contributes independent update to $w$.
@@ 3.2 Convergent Token-based Model Inversion (TMI-C)
TMI performs a single update per token per iteration. However, VLMs generate each token $y_{i}$ based on the preceding tokens $y_{<i}$. To better align with this generative dependency, we propose Convergent Token-based Model Inversion (TMI-C), which updates the latent vector $w$ multiple times for each target token before proceeding to the next. Specifically, for each token $y_{i}$, we perform $K$ updates to $w$, thereby encouraging convergence of the token-level inversion subproblem before advancing to $y_{i+1}$ (see Figure 1 (A) TMI-C). The details are presented in Algorithm 2.
@@
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},\dots,y_{m}),N,\lambda$
@@
4:for $i=1$ to $m$ do
5:  for $k=1$ to $K$ do
6:   Compute $\mathcal{L}$ using Eqn. (2).
7:   $w=w-\lambda\frac{\partial\mathcal{L}}{\partial w}$
8:  end for
@@ 3.3 Sequence-based Model Inversion (SMI)
Token-based model inversion methods treat each token independently, optimizing the latent vector $w$ based on individual token-level losses. As the output of VLMs is a sequence of tokens, we propose Sequence-based Model Inversion (SMI), which performs a single gradient update to $w$ by averaging the loss across all $m$ tokens in the sequence (see Figure 1 (A) SMI). By aggregating token-level losses into a unified objective, SMI leverages the interdependencies among tokens and provides more coherent gradients that reflects the structure of the full sequence. This global view encourages the model to recover a latent representation that is consistent across the entire sequence, rather than optimizing for each token in isolation. The details are presented in Algorithm 3.
@@
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},\dots,y_{m}),N,\lambda$
@@
3:for $k=1$ to NN do
4:   $\hskip-28.45274pt\mathcal{L}=\frac{1}{m}\sum_{i=1}^{m}\mathcal{L}_{inv}(M(\mathbf{t},G(w),y_{<i}),y_{i})$ (3)
5:  $w=w-\lambda\frac{\partial\mathcal{L}}{\partial w}$
6:end for
4 Sequence-based Model Inversion with Adaptive Token Weighting (SMI-AW)
In this section, we further propose a novel sequence-based MI with dynamic weighting. VLMs generate each output token $y_{i}$ based on the preceding text tokens $y_{<i}$ and the image $\mathbf{x}$. We observe that different output tokens have varying degrees of dependence on the visual input — some are strongly visually grounded, while others are less visually grounded and they are driven by prior linguistic context instead (Figure 1 (B), Figure 2). Consequently, the gradients of output token $y_{i}$ vary in informativeness for reconstructing images during MI.
If a token $y_{i}$ exhibits strong visual attention, it is likely more visually dependent, and its loss gradient carries richer visual information about the image. In other words, the strength of cross-attention could indicate how sensitive the token’s prediction is to the image content, which directly determines how informative its gradient is for model inversion. Therefore, we propose to use the magnitude of the attention map as a proxy for the informativeness of each token’s loss gradient in a model inversion step and use it to weight its contribution to the overall inversion gradient — tokens with higher visual attention receive larger weights, while those with weaker visual grounding are down-weighted. Note that the magnitude of the attention map can be readily obtained in white-box MI.
Let $\alpha_{i}$ denote the total visual attention value of the output token $y_{i}$. The corresponding weight $\beta_{i}$ for each output token $y_{i}$ is then computed as:
$\beta_{i}=\frac{\alpha_{i}}{\sum_{j=1}^{m}\alpha_{j}}$ (4)
Furthermore, we update these weights $\beta_{i}$ dynamically across inversion steps, since a token’s dependence on visual input can change as the reconstructed image gradually becomes more consistent with the target output. The method is presented in Algorithm 4. Overall, this adaptive weighting enables the optimization to focus on visually-grounded output tokens, producing gradients that more effectively guide the reconstruction of the private training image.
Figure 2: Analysis of visual–textual attention across output tokens and inversion steps. We visualize the cross-attention map between the reconstructed image and each output token during inversion. Different tokens exhibit markedly different attention maps: visually grounded tokens show strong attention, while others produce weak responses, indicating limited reliance on the image. Moreover, attention patterns evolve over inversion steps, as a token’s dependence on visual input changes when the reconstructed image becomes more consistent with the target output. These observations reveal that token-level gradients vary substantially in visual informativeness both across tokens and over time. This motivates our SMI-AW method, which dynamically reweights token contributions based on their visual attention strength. Additional attention map analysis can be found in Supp.
1:Input: $M,G,\mathbf{t},\mathbf{y}=(y_{1},\dots,y_{m}),N,\lambda$
2:Output: $G(w)$
3:for $k=1$ to NN do
4:  Compute $\beta_{i}$ for each token $y_{i}$ using Eqn. (4)
5:   $\mathcal{L}=\sum_{i=1}^{m}\beta_{i}\mathcal{L}_{inv}(M(\mathbf{t},G(w),y_{<i}),y_{i})$ (5)
6:  $w=w-\lambda\frac{\partial\mathcal{L}}{\partial w}$